OpenStudy (anonymous):

e^4(-2log,x,8)=e^24 , ,=base

sam (.sam.):

e^(4(-2log^x(8)))=e^(24) e^(-8log^x(8))=e^(24) ln(e^(-8log^x(8)))=ln(e^(24)) -8log^x(8)ln(e)=ln(e^(24)) -8log^x(8)ln(e)=(24) log^x(8)=-(3)/(ln(e)) x^(log^x(8))=x^(-(3)/(ln(e))) 8=x^(-(3)/(1)) 8=(1)/(x^(3)) (1)/(x^(3))=8 (1)/(8)=x^(3) x^(3)=(1)/(8) \[x=\sqrt[3]{1/8}\]